Data for Barry Computer Co. and its industry averages follow. The firm's debt is priced at par, so the market value of its debt equals its book value. Since dollars are in thousands, number of shares are shown in thousands too.

Barry Computer Company:
Balance Sheet as of December 31, 2019 (In Thousands)
Cash $ 68,640 Accounts payable $ 145,860
Receivables 231,660 Other current liabilities 68,640
Inventories 205,920 Notes payable to bank 77,220
   Total current assets $ 506,220    Total current liabilities $ 291,720
Long-term debt 197,340
Net fixed assets 351,780 Common equity (36,894 shares) 368,940
Total assets $ 858,000 Total liabilities and equity $ 858,000
Barry Computer Company:
Income Statement for Year Ended December 31, 2019 (In Thousands)
Sales $ 1,300,000
Cost of goods sold
   Materials $559,000
   Labor 338,000
   Heat, light, and power 39,000
   Indirect labor 78,000
   Depreciation 65,000 1,079,000
Gross profit $ 221,000
Selling expenses 156,000
General and administrative expenses 13,000
   Earnings before interest and taxes (EBIT) $ 52,000
Interest expense 15,787
   Earnings before taxes (EBT) $ 36,213
Federal and state income taxes (25%) 9,053
Net income $ 27,160
Earnings per share $ 0.7362
Price per share on December 31, 2019 $ 13.00
  1. Calculate the indicated ratios for Barry. Do not round intermediate calculations. Round your answers to two decimal places.
    Ratio Barry              Industry Average
    Current × 1.75 ×
    Quick    × 0.98 ×
    Days sales outstandinga days 30    days
    Inventory turnover    × 6.89 ×
    Total assets turnover    × 1.71 ×
    Profit margin     % 1.95 %
    ROA      % 3.33 %
    ROE      % 7.83 %
    ROIC      % 7.40 %
    TIE   × 3.26 ×
    Debt/Total capital      % 41.37 %
    M/B    4.30
    P/E    20.06
    EV/EBITDA    8.62

    aCalculation is based on a 365-day year.
Answer #1

Calculation of the required ratios :

1. Current Ratio = Current Assets / Current liabilities = 506,220 / 291,720 = 1.73 times

2. Quick Ratio = Quick Assets (Current assets - Inventories) / Current liabilities = (506,220-205,920) / 291,720 = 1.03 times

3. Days sales outstanding = (Average Accounts Receivable / Total credit sales)*No. of days = (231,660/13,00,000) * 365 days = 65.04 days

4. Inventory Turnover Ratio = Cost of goods sold / Average Inventory = 10,79,000 / 205,920 = 5.24 times

5. Total assets turnover ratio = Net sales / Total assets = 13,00,000 / 858,000 = 1.52 times

6. Profit margin = Net Income / Net sales = 27,160 / 13,00,000 = 2.09 %

Note : Here we have used Net income for calculating profit margin. Students can give a note and use any other profit such as EBIT or Gross profit for calculation of the same.

7. ROA = Net income / Average Total assets = 27,160 / 858,000 = 3.17 %

8. ROE = Net income / Shareholders' Equity = 27,160 / 368,940 = 7.36 %

9. Return on Invested Capital = Net Operating Profit after tax / Total Invested Capital (Debt+Equity) = 27,160 / (368,940+197,340) = 4.8 %

10. Times Interest Earned Ratio (TIE) = EBIT / Interest expense = 52,000 / 15,787 = 3.29 times

11. Debt/Total Capital = Debt/(Debt + Equity) = 197,340 / (197,340+368,940) = 34.85 %

12. Market to Book Ratio (MB) = Total Market capitalization / Book value = No. of shares * MPS / Book Value = (36894*13) / 368940 = 1.3 times

13. PE Ratio = MPS/EPS = 13/0.7362 = 17.66 times

14. Calculation of EV/EBITDA Multiple :

Enterprise Value = Market Capitalization + Total value of Debt - Cash & Cash Equivalents = (36,894*13) + (291,720+197,340) - 68,640 = 900,042 $

EBITDA = EBIT + Depreciation = 52,000 + 65,000 = 117,000 $

Hence EV/EBITDA = 900,042/117,000 = 7.69 times